OpenStudy (anonymous):

Please explain! +1 Medal!! Solve w -4w + 5 / 3 = -13 A. w = -11 B. w = -4 1/2 C. w = 7 3/4 D. w = 11

OpenStudy (alexandervonhumboldt2):

-4w=-39/3-5/3 -4w=-44/3

OpenStudy (alexandervonhumboldt2):

w=44/3*1/4

OpenStudy (alexandervonhumboldt2):

w=11/3

OpenStudy (campbell_st):

is the question 1. \[\frac{-4w + 5}{3} = -13\] or 2. \[-4w + \frac{5}{3} = -13\]

OpenStudy (anonymous):

.1

OpenStudy (campbell_st):

ok.... multiply both sides of the equation by 3, with remove the denominator so -4w + 5 = -39 does that make sense..?

OpenStudy (anonymous):

yes, thanks.

OpenStudy (campbell_st):

can you solve it from here, or do you need further help..?

OpenStudy (campbell_st):

Oh, and the answer isn't the one that has been posted

OpenStudy (anonymous):

I know, and yeah, i still need help haha

OpenStudy (campbell_st):

ok... so next subtract 5 from both sides of the equation. What do you get..?

OpenStudy (anonymous):

-1w = -34 ....?

OpenStudy (campbell_st):

no -4w + 5 = -39 -5 -5 -------------- -4w = -44 last step is divide both sides by -4

OpenStudy (anonymous):

ohhhh 11 i was overthinking it! Thanks @campbell_st

OpenStudy (campbell_st):

here is an alternate method called backtracking .... it uses a road map read the equation start with w multiply it by -4 then add 5 divide everything by 3 it all equations -13 '